1kg 4 kg 60°

A 4 kg mass and a 1 kg mass are suspended from a frictionless, 200g pulley of radius 10 cm, over a 60° incline as shown below. (friction on slope: µs=0.4, µk=0.1) What is the rate of acceleration of the 4kg mass (m/s2)?
Answer #1

Ansi FBD. To tal 1 600 0. tonig Top mgsino Hking coso Eqoutcome mig -Tisno o Ti-Tale=1 / Meyo 10 T1-72= ma 272 - Magsing not Mk m2gCOS® - Mza Adding ears cll (21 8 (32 M2gsinot ik m2 grose (mitm nith tma t mig 2a 1x9.81-4X981x sin60+X 4X981 Com lt 4+ 0.2 2 a=-4-355 M/S². negative sign implics, acceleration is downwards to slope. a = 4